OpenStudy (help_people):

Solve 3x2 + 18x + 15 = 0 by completing the square. 3(x + 3)2 = 6; x = −3 + square root 2, x = −3 − square root 2 3(x + 9)2 = 12; x = −11, x = −7 3(x + 3)2 = 12; x = −5, x = −1 3(x + 9)2 = 228; x = −9 + square root 76, x = −9 − square root 76

OpenStudy (anonymous):

divide by 3

OpenStudy (anonymous):

\[x^2+6x+5=0\] \[x^2+6x=-5\] add both sides \[\left( \frac{ 6 }{ 2 } \right)^2 ~i.e.,9\] \[x^2+6x+9=-5+9\] \[\left( x+3 \right)^2=2^2\] \[x+3=\pm2,x=?\]

OpenStudy (help_people):

4?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x+3=2 x=?

OpenStudy (help_people):

-1

OpenStudy (anonymous):

correct and x+3=-2 x=?

OpenStudy (help_people):

-5

OpenStudy (anonymous):

correct these are two values of x

OpenStudy (help_people):

THANK YOUUUU :)

OpenStudy (anonymous):

yw